1.     Two forces act on a parachutist during a drop. One force is the weight W(500N) of the parachutist. The other force is the upwards pull P(400N) from the parachute harness on the parachutist. The parachutist has a mass of 50 kg.


a.     What is the size of the unbalanced force acting on the parachutist?

b.     What is the acceleration of the parachutist?


1.     A boat engine produces a  force of 10000 N and the friction and water resistance total 3500 N.  If the mass of the boat is 2000 kg, what will be its acceleration?


1
Expert's answer
2021-05-17T13:31:51-0400

"1."

"\vec F_1=500;\vec F_2=400;m = 50"

"a)\vec F =\vec F_1+\vec F_2=500-400 =100"

"b)\vec F= m\vec a"

"\vec a = \frac{\vec F}{m}=\frac{100}{50}=2"

"\text {Answer:}\vec F=100N;\vec a = 2\frac{m}{s^2};"

"2."

"\vec F_1=10000;\vec F_2=3500;m = 2000"

"\vec F =\vec F_1+\vec F_2=10000-3500=6500"

"\vec F= m\vec a"

"\vec a = \frac{\vec F}{m}=\frac{6500}{2000}=3.25"

"\text {Answer: }\vec a = 3.25\frac{m}{s^2};"